OpenStudy (anonymous):

what percent of 67 is 33

OpenStudy (tkhunny):

100% of 67 is 67 50% of 67 is 33.5 <== Wow! That was close.

OpenStudy (kkutie7):

33/100=67/x 33x=6700 x= 203.03 this is an estimate. 203.03*.33=69.99

OpenStudy (anonymous):

wait so its __% of 67 is 33% right?

OpenStudy (tkhunny):

The point of my little demonstration was to FIRST give a clue what the right answer MIGHT be. Since we now know that the right answer is just under 50%, any other solution would have to be looked upon with disdain.